# Collapsing Domain Wall Networks: Impact on Pulsar Timing Arrays and Primordial Black Holes

### v1

@@ Abstract
Unstable domain wall (DW) networks in the early universe are cosmologically viable and can emit a large amount of gravitational waves (GW) before annihilating. As such, they provide an interpretation for the recent signal reported by Pulsar Timing Array (PTA) collaborations. A related important question is whether such a scenario also leads to significant production of Primordial Black Holes (PBH). We investigate both GW and PBH production using 3D numerical simulations in an expanding background, with box sizes up to $N=3240$, including the annihilation phase. We find that: i) the network decays exponentially, i.e. the false vacuum volume drops as $\sim\exp(-\eta^{3})$, with $\eta$ the conformal time; ii) the GW spectrum is larger than traditional estimates by more than one order of magnitude, due to a delay between DW annihilation and the sourcing of GWs. We then present a novel semi-analytical method to estimate the PBH abundances: rare false vacuum pockets of super-Hubble size collapse to PBHs if their energy density becomes comparable to the background when they cross the Hubble scale. Smaller (but more abundant) pockets will instead collapse only if they are close to spherical. This introduces very large uncertainties in the final PBH abundance. The first phenomenological implication is that within these uncertainties it is not possible to rule out the DW interpretation of the PTA signal. Second, in a different parameter region, there is the interesting possibility of producing all of the dark matter in the form of asteroid-mass PBHs from the DW collapse. Remarkably, this would also lead to a GW background in the observable range of LIGO-Virgo-KAGRA and future interferometers, i.e. Einstein Telescope and LISA.
@@ I Introduction
Overall, these aspects motivate a detailed investigation of the evolution of a DW network during its collapse phase (see [18, 19, 20, 21, 22, 23] for previous work) and of its gravitational relics, i.e. GWs (see [24, 25, 26, 27]) and PBHs, which we aim to perform in this work. We consider the simple case of DWs with a double-well potential, with an energy per unit area (tension) $\sigma$, and where the potential is slightly tilted by a term of size $\Delta V$, i.e. a bias such that one of the $\mathbb{Z}_{2}$-symmetric minima becomes a false vacuum [28] (see e.g. [29, 30, 31, 32] for other mechanisms to have viable long-lived walls). We simulate the corresponding DW network throughout the formation, scaling and annihilation regimes, using field theory numerical simulations in an expanding radiation-dominated background, with box size up to $3240^{3}$, while computing the GWs radiated throughout the evolution.
@@
Our numerical results will also allow us to establish the time evolution of the decaying network, in particular of its FV pockets. Indeed, the collapse of the network takes some time: the abundance of Hubble-sized FV pockets at some point drops very quickly, but a small fraction of rare super-Hubble sized pockets survive for a longer time as they must cross the Hubble radius to annihilate. Their abundance dramatically decreases in time, but their likelihood to collapse into a BH grows instead simply because the Schwarzschild radius associated with the FV pocket grows faster in time than the Hubble length. This may result in a tiny population of BHs at formation, but potentially large at present if the network collapses in the early universe (this formation mechanism is similar to the one in [34, 35] for isolated DWs, with the crucial difference that for a network the collapse is in general far from spherical).
@@
II Domain Wall Networks:
Scaling and Annihilation regimes
@@ III.4 Gravitational Waves
A simple estimate of the maximal energy density fraction (i.e. the energy density in GWs compared to that of the radiation background) is provided by the quadrupole formula, which gives $\Omega_{\text{gw}}(\eta)\sim 3/(32\pi)\alpha^{2}_{\text{tot}}(\eta)$, where $\alpha_{\text{tot}}\equiv\rho_{\phi}/(3H^{2}M_{p}^{2})$. This would correspond to the case in which all the energy density in the scalar field sources GWs. We compare our results with this simple estimate in Fig. 8, for three different choices of bias size for which we are able to capture the full GW production (with same initial conditions as in all previous figures). For all our choices, we find that the GW energy density fraction reaches a peak at $\eta\gtrsim 40$. The efficiency factor with respect to the simple quadrupole estimate at peak production is $\epsilon\equiv\Omega_{\text{gw}}/(3/(32\pi)\alpha^{2}_{\text{tot}})\simeq 0.5-0.6$. Fig. 9 shows that GWs are compatible with being mostly sourced by the kinetic component of the energy density in the scalar field, with a subdominant contribution from the gradient component, for one example value of bias size.
@@
Physically, our findings suggest that efficient GW production continues throughout the annihilation epoch, and so beyond the production during scaling studied in previous works, and that most of the relic abundance of GWs is determined by the late stages of the DW network collapse.
@@ V Phenomenological Implications
We now proceed to discuss the phenomenological impact of our estimates of the spectrum of stochastic GWs and of the abundance of PBHs from the DW network. We base our results on the numerical output described in Sec. III and on the analytical understanding of the evolution of the false vacuum fraction described in Sec. IV. We start by estimating the minimal PBH abundance, formed from Hubble sized PBHs pockets that reach $\alpha_{\text{loc}}=1$, using eq. (11). The total abundance may be expressed as a function of the energy fraction of the network at the time of GW emission, $\alpha_{\text{gw}}$, and the background temperature at that time, $T_{\text{gw}}$. The mass of these black holes is set roughly by the total energy in the Hubble volume and more precisely by eq. (22). The abundance at a given mass is experimentally constrained by a wide variety of probes. In Fig. 12 we translate we translate the bounds on the PBH abundance from [46] and [47] into bounds on the $\alpha_{\text{gw}}-T_{\text{gw}}$ plane (thick blue curve). We relate $\alpha_{\text{loc}}$ with $\alpha_{\text{gw}}$ using eq. (23), which is a minor refinement of eq. (16). Constraint on $\alpha_{\text{gw}}$ are obviously stronger for larger $T_{\text{gw}}$, as PBHs redshift as matter for a longer time. As an interesting example of the typical mass of these PBHs, we show in green the boundaries of the asteroid mass range $10^{-16}M_{\odot}\lesssim M_{\text{PBH}}\lesssim 10^{-11}M_{\odot}$ where the PBHs can account for the whole of the dark matter.
@@
In the same plot, we also show the range of parameters where the GW signal from the DW network could be observed by different GW detectors, from SKA [48] at the lowest frequencies to LIGO-Virgo-KAGRA (LVK) [36], ET and CE at the highest ones. We have fixed the frequency of the GW spectrum at the peak to be dictated by the Hubble radius at $T_{\text{gw}}$, and amplitude given by eq. (4), with efficiency $\epsilon=0.6$, as obtained from the numerical results, and assumed the spectrum to decrease as $1/\omega$ for frequencies $\omega$ larger than the peak and $\omega^{3}$ for smaller frequencies, based on [26]. We then plotted the regions of parameters where the spectrum overlaps with the power law sensitivity curves derived in [49]. In the same figure, we also show the current bounds obtained in [5] with LIGO-Virgo O3 data (LV) and the region of parameter space which could provide an interpretation of the PTA signal (red contours) [11].
Let us focus on the PTA region first. The bounds on PBHs formed from DWs with $\alpha_{c}=1$ do not exclude the DW interpretation of the signal. However, we note that the sub-Horizon population could potentially be in tension with the data if the threshold for PBH formation were as low as $\alpha_{c}=0.1$. Therefore, we conclude that the DW hypothesis remains safely open given the very large uncertainties, and more work is needed to better characterize this sub-Hubble population. Note also that our results differ drastically both from [17] (for example regarding the time dependence of the DW decay) and from the earlier contradictory claim [50].
As mentioned above, a particularly interesting region of parameter space is the asteroid mass range, where the totality of dark matter could be explained by PBHs from the network. This mass range is typically very hard to probe, given the particle-like size of their Schwarzschild radius. Crucially however, if PBHs originate from the DW network, a complementary GW signature of their existence is expected, that can be probed partially by LVK at design sensitivity and fully by ET and CE.
@@
Figure 12: Constraints on the PBH abundance from the collapse of the DW network for different values of the collapse threshold $\alpha_{c}$ (blue lines) in terms of the fraction of the energy density in the DW network at the time of GW emission ($\alpha_{\text{gw}}$) and the background temperature at that point ($T_{\text{gw}}$). The red ellipsis shows the region of parameters where the DW network interpretation explains the recent GW signals detected at PTA [11] at one and two standard deviations. The shaded regions are the constraints on the GW spectrum from LIGO-VIRGO O3 data (LV) and the prospects for detection with LIGO-Virgo-KAGRA design sensitivity (LVK), Einstein Telescope, Cosmic Explorer, LISA and SKA using the power-law integrated sensitivity curves from [49]. Finally, the green band corresponds the asteroid mass range ($10^{-16}-10^{-11}\,M_{\odot}$) for $\alpha_{c}=1$. For the other values of $\alpha_{c}$ the band moves slightly to the left.
@@ Acknowledgements.
We thank Daniel Figueroa and Francisco Torrentí for help with Cosmolattice. The work of F.R. is partly supported by the grant RYC2021-031105-I from the Ministerio de Ciencia e Innovación (Spain). F.R. thanks the MITP in Mainz (DE) for kind hospitality during completion of this work. R.Z.F. acknowledges the financial support provided by FCT – Fundação para a Ciência e Tecnologia, I.P., reference 2022.03283.CEECIND as well as the FCT projects CERN/FIS-PAR/0027/2021, UIDP/04564/2020 and UIDB/04564/2020. The work of A.N. is supported by the grants PID2019-108122GB-C32 from the Spanish Ministry of Science and Innovation, Unit of Excellence María de Maeztu 2020-2023 of ICCUB (CEX2019-000918-M) and AGAUR2017-SGR-754. A.N. is grateful to the Physics Department of the University of Florence for the hospitality during the course of this work. OP acknowledge the support from the Departament de Recerca i Universitats from Generalitat de Catalunya to the Grup de Recerca ‘Grup de Física Teòrica UAB/IFAE’ (Codi: 2021 SGR 00649) and the Spanish Ministry of Science and Innovation (PID2020-115845GB-I00/AEI/10.13039/501100011033). IFAE is partially funded by the CERCA program of the Generalitat de Catalunya.
@@ Appendix B FV pockets in the Nambu Goto approximation
Figure 13: The (conformal) time of collapse, $\Delta\eta$, of super-Hubble FV pockets depends mildly on the shape and it scales linearly with the initial size $R_{0}$. In blue, the time to reach $R(\eta)=0$, in red, time to enter the Hubble radius, $R(\eta)=\eta$.
@@
One expects a ‘burst’ of GWs from the collapsing FV regions as they shrink because they are significantly non-spherical. This GW production time $\eta_{\text{gw}}$ is then expected to be near the collapse time because it is dominated by the most numerous pockets, whith sizes of order $\eta_{\Delta V}$. In summary, the fact that Hubble-sized structures have to reach small sizes naturally leads to the expectation

### v3

@@ Abstract
Unstable domain wall (DW) networks in the early universe are cosmologically viable and can emit a large amount of gravitational waves (GW) before annihilating. As such, they provide an interpretation for the recent signal reported by Pulsar Timing Array (PTA) collaborations. A related important question is whether such a scenario also leads to significant production of Primordial Black Holes (PBH). We investigate both GW and PBH production using 3D numerical simulations in an expanding background, with box sizes up to $N=3240$, including the annihilation phase. We find that: i) the network decays exponentially, i.e. the false vacuum volume drops as $\sim\exp(-\eta^{3})$, with $\eta$ the conformal time; ii) the GW spectrum is larger than traditional estimates by more than one order of magnitude, due to a delay between DW annihilation and the sourcing of GWs. We then present a novel semi-analytical method to estimate the PBH abundances: rare false vacuum pockets of super-Hubble size collapse to PBHs if their energy density becomes comparable to the background when they cross the Hubble scale. Smaller (but more abundant) pockets will instead collapse only if they are close to spherical. This introduces very large uncertainties in the final PBH abundance. The first phenomenological implication is that the DW interpretation of the PTA signal is compatible with observational constraints on PBHs, within the uncertainties. Second, in a different parameter region, the dark matter can be entirely in the form of asteroid-mass PBHs from the DW collapse. Remarkably, this would also lead to a GW background in the observable range of LIGO-Virgo-KAGRA and future interferometers, such as LISA and Einstein Telescope.
CERN-TH-2024-020
@@ I Introduction
Overall, these aspects motivate a detailed investigation of the evolution of a DW network during its collapse phase (see [18, 19, 20, 21, 22, 23] for previous work) and of its gravitational relics, i.e. GWs (see [24, 25, 26, 27, 28, 29]) and PBHs, which we aim to perform in this work. We consider the simple case of DWs with a double-well potential, with an energy per unit area (tension) $\sigma$, and where the potential is slightly tilted by a term of size $\Delta V$, i.e. a bias such that one of the $\mathbb{Z}_{2}$-symmetric minima becomes a false vacuum [30] (see e.g. [31, 32, 33, 34] for other mechanisms to have viable long-lived walls). We simulate the corresponding DW network throughout the formation, scaling and annihilation regimes, using field theory numerical simulations in an expanding radiation-dominated background, with box size up to $3240^{3}$, while computing the GWs radiated throughout the evolution.
@@
Our numerical results will also allow us to establish the time evolution of the decaying network, in particular of its FV pockets. Indeed, the collapse of the network takes some time: the abundance of Hubble-sized FV pockets at some point drops very quickly, but a small fraction of rare super-Hubble sized pockets survive for a longer time as they must cross the Hubble radius to annihilate. Their abundance dramatically decreases in time, but their likelihood to collapse into a BH grows instead simply because the Schwarzschild radius associated with the FV pocket grows faster in time than the Hubble length. This may result in a tiny population of BHs at formation, but potentially large at present if the network collapses in the early universe (this formation mechanism is similar to the one in [36, 37] for isolated DWs, see also [38], with the crucial difference that for a network the collapse is in general far from spherical).
@@
II Domain Wall Networks: Scaling and Annihilation regimes
@@ III.4 Gravitational Waves
A simple estimate of the maximal energy density fraction (i.e. the energy density in GWs compared to that of the radiation background) is provided by the quadrupole formula, which gives $\Omega_{\text{gw}}(\eta)\sim 3/(32\pi)\alpha^{2}_{\text{tot}}(\eta)$, where $\alpha_{\text{tot}}\equiv\rho_{\phi}/(3H^{2}M_{p}^{2})$. This would correspond to the case in which all the energy density in the scalar field sources GWs. We compare our results with this simple estimate in Fig. 8, for three different choices of bias size for which we are able to capture the full GW production (with same initial conditions as in all previous figures). For all our choices, we find that the GW energy density fraction reaches a peak at $\eta\gtrsim 40$. The efficiency factor with respect to the simple quadrupole estimate at peak production is $\epsilon\equiv\Omega_{\text{gw}}/(3/(32\pi)\alpha^{2}_{\text{tot}})\simeq 0.5-0.6$. Fig. 9 shows that GWs are compatible with being mostly sourced by the kinetic component of the energy density in the scalar field, with a subdominant contribution from the gradient component, for one example value of bias size. The GW energy density fraction is computed by integrating the numerically obtained GW spectrum, although the final result is dominated by the region around the peak of the spectrum.
@@
Physically, our findings suggest that efficient GW production continues throughout the annihilation epoch, and so beyond the production during scaling studied in previous works, and that most of the relic abundance of GWs is determined by the late stages of the DW network collapse. Compared to the previous literature, the delay amounts to a factor of ${}1.3\times 1.4\sim 1.8$ in the time of GW emission, which was previously estimated to be $\eta_{\Delta V}$, and so, following eq. (4), an enhancement of $1.8^{4}\sim 10$ in the total GW abundance.
@@ V Phenomenological Implications
We now proceed to discuss the phenomenological impact of our estimates of the spectrum of stochastic GWs and of the abundance of PBHs from the DW network. We base our results on the numerical output described in Sec. III and on the analytical understanding of the evolution of the false vacuum fraction described in Sec. IV. We start by estimating the minimal PBH abundance, formed from Hubble sized PBHs pockets that reach $\alpha_{\text{loc}}=1$, using eq. (11). The total abundance may be expressed as a function of the energy fraction of the network at the time of GW emission, $\alpha_{\text{gw}}$, and the background temperature at that time, $T_{\text{gw}}$. The mass of these black holes is set roughly by the total energy in the Hubble volume and more precisely by eq. (22). The abundance at a given mass is experimentally constrained by a wide variety of probes. In Fig. 12 we translate the bounds on the PBH abundance from [49] and [50] into bounds on the $\alpha_{\text{gw}}-T_{\text{gw}}$ plane (thick blue curve). We relate $\alpha_{\text{loc}}$ with $\alpha_{\text{gw}}$ using eq. (23), which is a minor refinement of eq. (16). Constraint on $\alpha_{\text{gw}}$ are obviously stronger for larger $T_{\text{gw}}$, as PBHs redshift as matter for a longer time. As an interesting example of the typical mass of these PBHs, we show in green the boundaries of the asteroid mass range $10^{-16}M_{\odot}\lesssim M_{\text{PBH}}\lesssim 10^{-11}M_{\odot}$ where the PBHs can account for the whole of the dark matter [51].
@@
In the same plot, we also show the range of parameters where the GW signal from the DW network could be observed by different GW detectors, from SKA [52] at the lowest frequencies to LIGO-Virgo-KAGRA (LVK) [39], ET and CE at the highest ones. We have fixed the frequency of the GW spectrum at the peak to be dictated by the Hubble radius at $T_{\text{gw}}$, and amplitude given by eq. (4), with efficiency $\epsilon=0.6$, as obtained from the numerical results, and assumed the spectrum to decrease as $1/\omega$ for frequencies $\omega$ larger than the peak and $\omega^{3}$ for smaller frequencies. This behavior corresponds to that observed in the scaling regime [26], and has been roughly confirmed recently also during the annihilation phase by [28]. The GW spectra computed in our simulations roughly agree with those works, although a dedicated study is necessary to firmly establish the high frequency slope.
We then plotted the regions of parameters where the spectrum overlaps with the power law sensitivity curves derived in [53]. In the same figure, we also show the current bounds obtained in [5] with LIGO-Virgo O3 data (LV), which already indirectly constraints the maximal PBH abundance for $\alpha_{c}=1$, and the region of parameter space which could provide an interpretation of the PTA signal (red contours) [11].
Let us focus on the PTA region first, i.e. at $T_{\text{gw}}\approx 1-10$ GeV. We find the DW interpretation of the PTA signal to be overall compatible with constraints on PBHs for collapse thresholds $\alpha_{c}\gtrsim 0.1$. Interestingly, if $\alpha_{c}\lesssim 1$, a fraction of PBH dark matter from DW collapse is expected, which is compatible with what is inferred from the BH merger rate measured by LIGO/Virgo. Significant tension between PTA observations and astrophysical bounds on PBHs would instead arise only if the collapse threshold were even smaller, i.e. $\alpha_{c}\lesssim 0.1$. The likelihood of such low thresholds remains to be assessed by future work. Our results differ drastically both from [17] (for example regarding the time dependence of the DW decay) and from the earlier contradictory in the first version of [54].
As mentioned above, a particularly interesting region of parameter space is the asteroid mass range, for $10^{6}\text{ GeV}\lesssim T_{\text{gw}}\lesssim 10^{9}\text{ GeV}$, where the totality of dark matter could be explained by PBHs from the network. This mass range is typically very hard to probe, given the particle-like size of their Schwarzschild radius. Crucially however, if PBHs originate from the DW network, a complementary GW signature of their existence is expected, that can be probed partially by LVK at design sensitivity and fully by ET and CE.
@@
The interplay of GW and PBHs signatures described in this work is similar to the more studied scenario of PBH formation from the collapse of large adiabatic perturbations from inflation. However, GWs from density perturbations as an interpretation of PTA data have been shown to be in tension with constraints from PBH overproduction [55, 56]. In contrast, the mechanism presented here remains viable according to our current understanding. Additionally, asteroid-mass PBH DM from inflationary perturbations has been argued to give a GW signal peaked in the LISA frequency band [57], whereas ground-based interferometers (LVK, ET, CE) are best-suited to indirectly probe PBH DM from DW collapse.
Figure 12: Constraints on the PBH abundance from the collapse of the DW network for different values of the collapse threshold $\alpha_{c}$ (blue lines) in terms of the fraction of the energy density in the DW network at the time of GW emission ($\alpha_{\text{gw}}$) and the background temperature at that point ($T_{\text{gw}}$). The region above the solid blue curve is conservatively excluded. The red ellipsis shows the region of parameters where the DW network interpretation explains the recent GW signals detected at PTA [11] at one and two standard deviations. The shaded regions are the constraints on the GW spectrum from LIGO-VIRGO O3 data (LV) and the prospects for detection with LIGO-Virgo-KAGRA design sensitivity (LVK), Einstein Telescope, Cosmic Explorer, LISA and SKA using the power-law integrated sensitivity curves from [53]. Finally, the green band corresponds the asteroid mass range ($10^{-16}-10^{-11}\,M_{\odot}$[58, 59] and the black dashed lines correspond to PBHs between 1 and 100 solar masses, for $\alpha_{c}=1$. For the other values of $\alpha_{c}$ the band moves slightly to the left.
@@ Acknowledgements.
We thank Daniel Figueroa and Francisco Torrentí for help with Cosmolattice. We thank Juan Sebastían Valbuena-Bermúdez for spotting some typos in the first version of this paper. The work of F.R. is partly supported by the grant RYC2021-031105-I from the Ministerio de Ciencia e Innovación (Spain). F.R. thanks the MITP in Mainz (DE) for kind hospitality during completion of this work. R.Z.F. acknowledges the financial support provided by FCT – Fundação para a Ciência e Tecnologia, I.P., reference 2022.03283.CEECIND as well as the FCT projects CERN/FIS-PAR/0027/2021, UIDP/04564/2020 and UIDB/04564/2020. The work of A.N. is supported by the grants PID2019-108122GB-C32 from the Spanish Ministry of Science and Innovation, Unit of Excellence María de Maeztu 2020-2023 of ICCUB (CEX2019-000918-M) and AGAUR2017-SGR-754. A.N. is grateful to the Physics Department of the University of Florence for the hospitality during the course of this work. OP acknowledge the support from the Departament de Recerca i Universitats from Generalitat de Catalunya to the Grup de Recerca ‘Grup de Física Teòrica UAB/IFAE’ (Codi: 2021 SGR 00649) and the Spanish Ministry of Science and Innovation (PID2020-115845GB-I00/AEI/10.13039/501100011033). IFAE is partially funded by the CERCA program of the Generalitat de Catalunya.
@@ Appendix B FV pockets in the Nambu Goto approximation
Figure 13: Ratio of the initial radius $R_{0}$ and the (conformal) time to reach $R(\eta)=0$, $\Delta\eta$, of super-Hubble FV pockets. It depends mildly on the shape and it asymptotes to a constant value (dotted black curve).
@@
One expects a ‘burst’ of GWs from the collapsing FV regions as they shrink because they are significantly non-spherical. This GW production time $\eta_{\text{gw}}$ is then expected to be near the collapse time because it is dominated by the most numerous pockets, with sizes of order $\eta_{\Delta V}$. In summary, the fact that Hubble-sized structures have to reach small sizes naturally leads to the expectation